OpenStudy (anonymous):

The tickets for a high school basketball game are $2.50 for students and $4.00 for adults. At Tuesday night's game, 95 tickets were sold for a total of $290.00. How many student tickets were sold?

OpenStudy (anonymous):

S=# of students A=# of adults 2.50S + 4.00A= 290 ($2.50 times the number of students plus $4.00 times the number of adults equals the amount of money collected $290) S + A = 95 Number of students tickets plus number of adults tickets equals number of tickets sold (95) Subtract S from both sides of last equation to solve for A A = 95 - S Plug A into the first equation 2.5S + 4(95-S)=290 2.5S + 380 - 4S = 290 - 1.5 S + 380 = 290 -1.5 S = -90 S = 60 60 Student tickets were sold.

OpenStudy (anonymous):

Thanks:)